OpenStudy (anonymous):

Solve 4 + 3^5x =8 for x.

OpenStudy (anonymous):

\[4+3^{5x}=8\] subtract \(4\) to get \[3^{5x}=4\] solve for \(5x\) using change of base \[5x=\frac{\log(4)}{\log(3)}\]

OpenStudy (anonymous):

so \[x=\frac{\log(4)}{5\log(3)}\]

OpenStudy (anonymous):

Ohh, I get it now. Thank you so much!